Actions
17/4/2015 : Weekend Activity Review¶
- Review of actions from Monday's meeting
- Fence from Q9 down to main door should be in place
- Cranes should be in right place
- Overview of readiness list from MDR1 and MDR2 : http://micewww.pp.rl.ac.uk/projects/operations/wiki/Operations_Readiness_To_Do_List
- minor changes to checklists (Steve) : Done
- Work on being able to access archiver data from external sites (Paolo/Pierrick) : (Stands)
- Get input on run plan from Jaroslaw (Jaroslaw/Steve/Milorad) : (see run plan)
- Steve to ask Chris to insert a line in the shift documentation asking the shifter to make sure
no other run control instance is operating before starting a new one : Done. Should be new instructions in the CR - Milorad to check with Debbie/Rose what the appropriate process is to get swipe access, possibly using the new ISIS Visitor Database. Done. Only the ISIS Crew can grant this. I am talking with Alan Stevens about an acceptable access procedure
- Steve to update MDR2-CMP-01. This will be available but in the new version of the framework. Done
- Milorad to talk to Yordan about attendance on Friday for start of run
- General comments about control room activities arising from last weekend run
- Please be methodical when going through checklists and ensure that all steps are carried out. (All)
- Enhance written Elog error messages so that physics shifter has a better idea of what went wrong and whether it is serious. Include error
message in Elog entry (All) - for the moment we will keep the Run Summary Spreadsheet going in parallel with the CDB (to be reviewed every week)
Check CDB after each run - put in Chris shift checklist. The CDB viewer can be found here
- Readiness of the downstream PID detectors
- TOF Gain of FADC issue: tof.png
- Review of run plan for April 18/19 : http://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20150418
- Today's plans
- BLOC Training
- Trainees : Mark (
MPB but will be available later in the day), Durga, Ray, Kevin, Chris Heidt, Alan Bross (
MPB but may be available later in the day). - Control room training : 1:00 pm
- Hall Closure : 3:30 pm : Full PPS tutorial required.
17/4/2015 : Minutes and Action Items arising from Weekend Activity Review¶
Action Items¶
- Pierrick (please) update his Operations readiness list entry
- Milorad to check with John about locking the floor lifting jack away.
Readiness of downstream PID detectors¶
Ready.
TOF Gain or FADC issues¶
We plan to use the reference run on Friday afternoon to try to diagnose the issue with the TOF
gains. Note that this issue may degrade the timing resolution but not the spatial resolution so that
the STEP I runs can go ahead as in the run plan. Durga plans to look into this issue, first by comparing
KL PMT response with TOF response to try to track the point where the change was introduced.
Run plan¶
Please check the run plan for measurement configurations and for notes specific to this run
Updated by Boyd, Steven about 8 years ago ยท 9 revisions